I am working on mixed numbers and variables in fractions. The problem is Solve for x: (2 1/2)/24)=x/30
I'm not quite sure how to start. Do I multiply both sides by 24 to leave the mixed number by its self which would give me 2 1/2 = 3x/5 then multiply the both sides by 5 to get 25/2 = 3x then divide by 3 to get 25/6. Or do I change the mixed number to a decimal so it is 2.5/24 = x/30. I'm just really confused...

Found 2 solutions by checkley79, solver91311:
Answer by checkley79(3341)

(2 1/2)/24)=x/30
((5/2)/24)=X/30 INVERT THE DENOMINATOR (24) & MULTIPLY
(5/2)(24/1)=X/30
120/2=X/30 CROSS MULTIPLY.
2X=120*30
2X=3600
X=3600/2
X=1800 ANS.
PROOF:
(5/2)/24=1800/30
120/2=1800/30
2*1800=120*30
3600=3600

Answer by solver91311(24713)



Either process results in the same answer -- given CORRECT arithmetic. In your first method you also need to convert your mixed number to an improper fraction.



Multiply by the LHS denominator. (Note: 6 goes into 24 four times and into 30 five times)



Now convert to an improper fraction. 2 is four halves plus one-half is



Multiply by the reciprocal of the coefficient on



Then you can integer divide 25 by 8 to convert back to a mixed number if that happens to tickle your fancy.

If you do it by changing everything to decimals, you end up with which is the same thing as 3 and 1/8 or 25/8.

Now for the way I would have done it in the first place.

Step 1: Multiply the LHS by 1 in the form





Multiply both sides by 48



Reduce the fraction in the RHS



Multiply by the reciprocal of the coefficient in the LHS



John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism